### v6

# Linear-time classical approximate optimization of rugged-energy-landscape cubic-lattice classical spin glasses

Abstract
Rugged energy landscapes present computational difficulty for optimization. A quantum alleviation may be possible via tunneling. Classically, an alleviation may arise through optimizing over subsystems and concatenating the results, but this introduces error in the optimization of the full problem and thereby raises the question of whether such an approach actually has a scaling advantage over approximately optimizing the full system directly to the same error. Here we investigate this question in the setting of cubic-lattice classical Ising spin glasses, where recent theoretical and experimental developments open the possibility of showing quantum speedup with quantum annealing, and where classical time-complexity results remain absent. For the subsystem-based approach we introduce a very simple deterministic tensor-network heuristic that features linear time and space complexity for approximate optimization of the full problem. For the full-system approach we use simulated annealing and parallel tempering. For the most rugged instances generated with tile planting on system sizes up to 565656, we find that full-system simulated annealing and full-system parallel tempering display a slightly superlinear time complexity when targeting the same energy error achieved by the subsystem-based linear-time heuristic. We also find that the error of the latter heuristic monotonically decreases with increasing ruggedness over the higher end of the ruggedness spectrum. These results suggest that subsystem-based classical optimization heuristics should be taken into account when seeking to demonstrate quantum speedup on rugged energy landscapes. We discuss prospects for reducing the error of our heuristic, for adapting our heuristic to arbitrary graphs, and for low-power, accelerated implementations with photonic matrix-multiplication hardware.
I Introduction
Classical spin-glass models have both fundamental and applied significance, and the simplest possible class of such Hamiltonians is the Ising spin glass in zero field, given by
where , , and and are site indices. Computing the ground state of Eq. (1) is known to be NP-hard [1], and a more structured choice of allows it to encode problems in discrete combinatorial optimization [2] such that the Hamiltonian encodes a cost function to be minimized and the ground state corresponds to the optimal solution of the optimization problem. On a -dimensional lattice with only nearest-neighbor interactions, when follows a random distribution (typically bimodal or Gaussian) this model is known as the classical Edwards-Anderson Ising model with zero field and is thought to capture the fundamental physics of spin-glass materials [3]. Yet, its low-temperature phase remains incompletely understood despite decades of effort [4].
I.1 Quantum speedup for approximate optimization
While computing the ground states (i.e., exact optimization) of classical spin glasses is therefore of both fundamental and applied interest, computing their low-energy states (i.e., approximate optimization) can also be useful. The low-energy excited states of classical spin-glass Hamiltonians, for instance, can provide additional insights into the low-temperature phases of classical spin-glass materials. For combinatorial optimization, merely reduced-cost solutions can be of practical benefit, and what is often desired is a set of low-energy states that satisfy various constraints [5, 6]. Even in cases where the optimal solution is the goal, the outputs of approximate optimization algorithms could potentially serve as warm starts (i.e., initial conditions that result in a shorter runtime compared to random initial conditions) for exact optimization algorithms.
There is a long history of investigating the possibility of quantum advantage (i.e., some type of performance advantage of quantum algorithms over classical algorithms) for the optimization of classical spin glasses [7, 8]. A key performance metric is the time-to-solution (TTS). In this work we are concerned specifically with the type of quantum advantage that is termed quantum speedup [9], which means a superior (i.e., smaller) time complexity (i.e., scaling of the TTS as a function of system size ()) compared to the best-known time complexity of classical algorithms for the same computational task. An advantage in time complexity is sometimes referred to as a scaling advantage.
For exact optimization there have been a number of results showing the potential for quantum speedup [10, 11, 12, 13, 14], though an actual demonstration remains absent. Regarding quantum speedups in approximate optimization, Refs. [15, 16] theoretically show the potential for it; Ref. [17] claims an actual demonstration of quantum speedup of D-Wave quantum annealing by comparing its time complexity to that of parallel tempering with isoenergetic cluster moves (a type of Markov-Chain-Monte-Carlo algorithm) for the approximate optimization of a class of quasi-2D spin glasses, but the more recent work of Ref. [18] refutes this claim by showing that the Simulated Bifurcation Machine [19, 20], a classical heuristic inspired by nonlinear Hamiltonian dynamics, achieves better time complexity than quantum annealing in the same class of problems. Thus it is apparent that claims of quantum speedup for classical spin glass optimization should not be based on comparisons with Markov-Chain-Monte-Carlo algorithms alone.
In this work we are concerned specifically with the issue of quantum speedup for the computational task of approximate optimization of Eq. (1) on the simple cubic lattice. To the best of our knowledge, this issue has not yet been addressed in the literature: the time complexity of neither classical nor quantum algorithms has been established for this computational task. Yet, the state-of-the-art of D-Wave quantum annealing encompasses approximate optimization of such models: D-Wave has demonstrated the approximate optimization of a cubic-lattice classical Ising spin glass with over 5,000 spins with pure quantum annealing [21] and over 11,000 spins with hybrid quantum annealing [22]. Experimental investigation of the time complexity of quantum annealing for this computational task is therefore feasible. Further, recent theoretical results [23, 24, 25, 26, 27] suggest the possibility of modifying quantum annealing so as to achieve a substantial improvement in the time complexity for this computational task compared to ordinary quantum annealing. It is therefore timely to ask for data on the time complexity of classical heuristics for the approximate optimization of cubic-lattice versions of Eq. (1); here we address this by presenting and testing a new classical heuristic that is tailored to such models, and also by providing a comparison of that new heuristic with the Markov-Chain-Monte-Carlo algorithms simulated annealing (SA) and parallel tempering (PT). Our results reveal that the new heuristic at least sometimes has a scaling advantage over SA and PT for approximate optimization of cubic-lattice classical Ising spin glasses. Thus, any attempts to show quantum speedup on this class of problems must take this new heuristic, and potential improvements thereof, into account.
I.2 Energy-landscape ruggedness as a source of computational difficulty
State-of-the-art Markov-Chain-Monte-Carlo (MCMC) algorithms, such as SA and PT, are often considered to be the best general classical heuristics for approximate and exact optimization of classical spin-glass models [17] (but see Ref. [18]). Consequently, computational hardness has become synonymous in the spin-glass literature with MCMC hardness (e.g., Ref. [28]). For this reason, Ref. [29] suggests that spin-glass models that display a steep increase in the autocorrelation time of parallel-tempering Monte Carlo as the temperature is lowered are good places to look for quantum advantage in optimization. The optimization hardness of certain spin glass instances for MCMC algorithms is understood to be due to the rugged energy landscape (i.e., a plethora of local minima) in spin-configuration space [30, 31]. Moreover, machine-learning enhancements of Monte-Carlo methods have been shown to be ineffective at overcoming this source of computational hardness [32], and non-MCMC classical heuristics that model dynamical systems evolution, such as the Simulated Bifurcation Machine mentioned previously, also display a substantial performance drop with increasing landscape ruggedness [33, 34]. In contrast, it is thought that quantum algorithms (such as quantum annealing) may be able to (at least partially) overcome this source of hardness via quantum tunneling [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45].
In this work we make the simple observation that a potential classical way of partially overcoming the difficulty of approximately optimizing over rugged energy landscapes is to optimize over subsystems and simply concatenate the results to approximate the solution to the full problem; this is a linear time heuristic. The intuition here is that if much or all of the energy landscape is very rugged, then efficiently arriving at a low-energy solution in this manner will bypass much of the ruggedness. Indeed, our results in this work do show a scaling advantage for our subsystem optimization-based heuristic compared to full-system MCMC on extremely rugged energy landscapes when targeting a solution that is within a few percent relative energy error of the ground state. While approximately solving larger optimization problems via optimizing their subsystems is not a new idea, to our knowledge this idea has not yet been investigated specifically as a means of alleviating the computational difficulty presented by energy-landscape ruggedness.
I.3 A subsystem optimization-based heuristic for cubic-lattice spin glasses
Our subsystem-based classical heuristic relies on a tensor-network representation of the partition function, and is made possible through the following line of developments: First, the work in Refs. [46, 47, 48, 49] presented ways of using tensor-network algorithms for homogeneous classical lattice models. This included a method of sampling from the Boltzmann distribution of homogeneous, two-dimensional classical spin lattices [49]. Then, Ref. [50] showed that the partition function of an inhomogeneous classical spin lattice may be exactly represented by the contraction of a network of tensors where the geometry of the network reflects that of the Hamiltonian’s interaction graph; while an exact contraction of the full network has an exponential cost, the contraction of the full network may be approximated, via truncated matrix decompositions, in polynomial time. Ref. [51] then, analogously to the work in Ref. [49] for homogeneous systems, demonstrated how to use this idea to sample (via computation of conditional marginals) the low-temperature Boltzmann distribution of planar and quasi-planar spin glasses in polynomial time. Finally, Ref. [52] used a technically different type of approximate contraction from Ref. [51] to generate near-optimal solutions of spin glasses on periodic square and cubic lattices in quadratic time. In both of the latter two works, the time complexity is determined by the complexity of contracting the network and not the Monte-Carlo hardness (which is related to the ruggedness of the free-energy landscape in configuration space), and the heuristic is successful in generating near-optimal solutions for all tested problem instances. We note, however, that the data in Ref. [52] for cubic-lattice spin glasses only spans two sizes (444 and 666) due to the high absolute time cost.
Compared to the approach in Ref. [52], the new method we present here constitutes a more efficient way of using tensor-network representations of classical, cubic-lattice spin-glass partition functions for approximate optimization. In Refs. [51] and [52], the amount of the network that is utilized to compute conditional marginals for single spins grows with the number of spins () that are conditioned upon (though Ref. [52] always uses the full network for simplicity). This leads to a time complexity for computing a total spin configuration that is quadratic in the total number of spins (an exponential complexity is avoided by utilizing truncated matrix decompositions). The method in the present work instead has a linear time (and memory) complexity for short-range classical spin glasses, and also eliminates the need for matrix decompositions. These improvements result in a dramatic reduction in walltime, and can be understood via a very simple intuition about short-range-correlated spin glasses: bootstrapping of approximate local energy minimization should lead to approximate global energy minimization. By “short-range-correlated" we mean that the short-range Hamiltonian does not amount to an embedding of a long-range Hamiltonian (as in, for example, Ref. [53]). Therefore, in contrast to the method in Ref. [52], the one presented here is able to reach large system sizes on the cubic lattice and thereby provide substantial data to challenge the time complexity of quantum algorithms for the approximate optimization of rugged-energy-landscape classical spin glasses on the cubic lattice.
Also in contrast to the tensor-network method in Ref. [52], the absence of matrix decompositions from the new method in this work means that its computational cost is dominated by matrix multiplications and it can therefore be substantially sped up with GPUs or FPGAs. Further, due to the ubiquity of matrix multiplication in computing, specialized photonic hardware for it, which seeks to improve both power consumption and speed over conventional processors, is under active development [54, 55, 56, 57]. The new heuristic thereby serves, in principle, as the basis for a new type of specialized computing machine for obtaining near-optimal solutions of certain types of discrete combinatorial optimization problems.
The approximate local energy minimizations in the new method are done via exact contractions of tensor-network fragments of fixed size, which is why bootstrapping them over the system requires no matrix decompositions and has () time complexity. While such a local optimization method will be biased against spin configurations with large-scale structures, such as droplets, it is intuitive that not all low-energy configurations will have such structures. However, small perturbations to the bonds in classical cubic-lattice spin glasses can substantially alter the ground state [58, 59], so it is not a priori obvious how small of an average error such a bootstrapping heuristic will yield, nor is it obvious how much fluctuation will be in that error from instance to instance within a fixed hardness class. Our results from all of the levels of energy-landscape ruggedness that we test reveal an upper bound on the average relative energy error of and a variance that varies widely over the different instance classes. The average-error upper bound is substantially smaller than the smallest upper bound of that is theoretically proven (assuming PNP) for the error of non-heuristic polynomial-time approximation algorithms for Eq. (1) (see End Matter of Ref. [60]).
The new heuristic presented here shares the local-optimization spirit of the algorithm in Sec. III. of Ref. [61], but it is different in the following crucial ways: (1) the present algorithm aims at only approximate optimization instead of exact optimization, (2) the present algorithm has a guaranteed linear time complexity, and (3) the present algorithm consists almost purely of matrix multiplications (i.e., exact contractions of tensor networks).
I.4 Outline for the rest of the work
In Section II we provide details of our tensor-network heuristic. In Section III.1 we present results from our heuristic for the cubic-lattice model (the model that is used in recent D-Wave experiments [21, 22]). In Section III.2 we present results from our heuristic on two classes of the cubic-lattice tile-planted-solution model across their full range of energy landscape ruggedness, and we present results from simulated annealing and parallel tempering on the most rugged instances of that same model. We conclude in Section IV with a summary and outlook.
II Tensor-Network heuristic details
As explained in Refs. [51, 52, 62] both conditional and unconditional single-spin marginals for a classical spin glass in the canonical ensemble can be computed via the contraction of a tensor network wherein the indices of a given tensor correspond to single-spin configurations and the elements of the tensor correspond to the Boltzmann weights of the joint configurations of the spins at the tensor’s legs (e.g., a two-index tensor contains the Boltzmann weights for all the possible configurations of two spins); the geometry of the tensor network mirrors the geometry of the Hamiltonian’s interaction graph. The contraction of the network results in a multiplication of the local Boltzmann weights across the whole system that yields the (conditional or unconditional, depending on details of the network) marginal for the spin of interest. See Fig. 1 and Appendix A of Ref. [52] for further explanation. Computation of conditional single-spin marginals, and single-spin decimations according to those single-spin marginals, allows one to sample the Boltzmann distribution
where s is a spin configuration vector for the entire system. For computing the exact ground state, exact contractions of the network are required with the inverse temperature set to infinity. While infinite is not numerically accessible, and exact contractions of the entire network are too costly for large systems, the intuition behind the heuristic in Ref. [52] is that approximate contractions of the entire tensor network with sufficiently large should yield low-energy spin configurations. For further details we refer the reader to Ref. [52].
The algorithm in this work is a modification of the one in Ref. [52]: the marginal of a single spin is (approximately) computed by exactly contracting only a fragment of the network that is local to the spin of interest. Sequentially computing single-spin marginals with tensor-network fragments that overlap previously decimated spins results in a bootstrapping of approximate local energy minimization if is sufficiently large (see Fig. 2 for an illustration on the square lattice). The intuition behind this approach is that if the Hamiltonian is not long-range correlated, then such a bootstrapping should also result in an approximate global energy minimization. Such an outcome would be consistent with the finding in Ref. [52] that approximate optimization can be successfully accomplished for short-range-correlated Hamiltonians with approximate contractions of the full tensor network that use only small bond dimensions. In this work we implement the new algorithm with the Python libraries quimb [63] and cotengra [64] on an Apple M2 Ultra CPU with 16 performance cores, 8 efficiency cores, and 128 GB of RAM. We note that in two and higher dimensions this algorithm is straightforward to parallelize by using multiple fragments simultaneously, though we do not do so here. Appendix A contains details of the SA and PT implementations.
III Experiments
III.1 cubic-lattice model
This model is described by Eq. (1) with chosen from a uniform distribution over . We use systems with dimensions spins and periodic boundary conditions. The relative energy error from the ground state is given by
where is the Hamiltonian energy of the computed configuration and is the ground state energy. We compute this by using the value of the ground state energy density for the thermodynamic limit that is numerically estimated in Ref. [65].
We first test the solution quality of the algorithm as a function of fragment size: we apply the bootstrapping algorithm to ten instances of with cubic fragments of size spins with and ; with we find the computation time of exactly contracting a single fragment to be impractically long. The results are shown in Fig. 3; as expected from the results in Ref. [66], the error decreases monotonically with , so we use cubic fragments with for the rest of the computations.
For and (ten instances each), the energy error data is shown in Fig. 4. The minimum error over the tested values of is , and the error fluctuation is less than 1%. The error is non-monotonic in due to finite numerical precision and possibly also finite fragment size. The TTS data in Fig. 10 of Appendix B closely follows the theoretical expectation of linear scaling.
III.2 cubic-lattice tile planting
This model is characterized in Ref. [31]. The model contains multiple base classes of instances; the three known as , , and can be generated by the Python library called Chook [67]. Following Ref. [31] we generate two families of instances: gallus_26 (a mixture of and ), and gallus_46 (a mixture of and ). Both families are parameterized by , which denotes the fraction of the full-system Hamiltonian that belongs to . For all instances we enable the option in the Chook library to scramble the ground states with gauge transformations. The error is the defined the same as in Eq. (3), except now is given exactly by the Chook library for each generated instance.
Between and , the hardness of exact optimization with MCMC of gallus_26 and gallus_46 is shown in Ref. [31] to increase (roughly) monotonically with increasing . When the MCMC exact-optimization hardness of gallus_46 is equivalent to that of the model, and the MCMC exact-optimization hardness of gallus_26 is slightly higher. When , gallus_26 and gallus_46 are equivalent and have an MCMC exact-optimization hardness a few orders of magnitude greater than the model.
We first test the performance of our heuristic on instances from the class (i.e., ) with different system sizes. As with the model, we use cubic fragments of size 555 and ten problem instances at each value of . The energy error data is shown in Fig. 5 and Fig. 6. The error is non-monotonic in due to finite numerical precision and possibly also finite fragment size, but monotonically decreases with increasing system size. Fig. 11 in Appendix B confirms that the TTS data closely follows the theoretical expectation of linear scaling.
From Fig. 6 we estimate an upperbound on the asymptotic error for the linear-time heuristic of 4% for the class, and we independently optimize the same instances from with SA and PT to 4% error. Plotting the data for SA and PT on semilog plots vs. and several smaller powers of (not shown) does not reveal a straight line, so we conclude that the SA and PT scaling is most likely polynomial. Fitting to finds for SA (Fig. 7) and for PT (Fig. 8), compared to 1.0081 for the tensor-network heuristic (Fig. 11). We note that this is even though the data that is used to compute the scalings for SA and PT is from runs with instance-specific tuning of the hyperparameters (see Appendix A). Thus we conclude a slight scaling advantage for the subsystem-based tensor-network heuristic over full-system SA and PT.
We next test our heuristic on gallus_26 and gallus_46 with and . The results are displayed in Fig. 9. We find an (empirical) upper bound on the error of about 7.5%, which is almost a factor of two better than the upper bound of 11.8% [60] known for non-heuristic approximate optimization algorithms for Eq. (1) that operate in polynomial time. We note, however, that this upper bound for non-heuristic approximate optimization takes into account the worst possible cases over all possible geometries for Eq. (1). Comparing to it the upper bound of our heuristic that is specialized to the cubic lattice is therefore not a completely fair comparison, but it is still of interest since it shows that our heuristic at least beats that more general upper bound of non-heuristic approximate algorithms. Interestingly, though Ref. [31] confirms increasing landscape ruggedness (i.e., MCMC hardness for exact optimization) as increases between 0.8 and 1, our linear-time heuristic actually shows a monotonically decreasing error.
IV Summary and Outlook
The question of where quantum speedups lie is an active area of research. It is thought that optimization of classical spin glasses with rugged energy landscapes may be a good place to look for such speedups due to the potential of quantum tunneling. Establishing a quantum speedup for a specific problem requires a comparison with the best known scaling behavior of classical heuristics on that same problem. Markov-Chain-Monte-Carlo (MCMC) algorithms, such as simulated annealing (SA) and parallel tempering (PT), have for a long time been considered the gold standard of classical heuristics for optimization of classical spin glasses with rugged energy landscapes. In this work we showed that a certain non-MCMC classical heuristic can sometimes have superior scaling to SA and PT for approximate optimization of rugged-energy-landscape classical Ising spin glasses on the cubic lattice. This suggests that classical heuristics aside from SA and PT should be taken into account when attempting to establish a quantum speedup on this type of problem, which is relevant to the state-of-the-art of D-Wave’s experimental capability for approximate optimization of classical spin glasses.
The non-MCMC classical heuristic that we introduced is actually one version of a meta-heuristic: instead of optimizing the full problem directly, it concatenates (via a bootstrapping procedure) the results of subsystem optimizations, but the subsystem optimizations can be done, in principle, with any heuristic. We used a tensor-network heuristic for the subsystem optimizations. The linear scaling of the meta-heuristic would not change by using a different heuristic for the subsystem optimizations, but there could be a big difference in the wall time. In principle, the tensor-network subsystem heuristic could outperform (both in terms of wall time and optimization error) other classical heuristics for the subsystem optimizations if the energy landscape of the subsystems was sufficiently rugged because the tensor-network subsystem heuristic does not perform a search over the energy landscape. Thus, the present classical version of the meta-heuristic may be a way to partially overcome the computational hardness of extremely rugged energy landscapes. On the other hand, quantum heuristics that exploit tunneling may also outperform search-based classical heuristics for optimization of subsystems with sufficiently rugged energy landscapes. All-in-all, certain versions of the present meta-heuristic have the potential to alleviate the computational hardness of rugged energy landscapes through two possible ways: 1) replacing full-system optimization with multiple subsystem optimizations, and 2) performing the subsystem optimizations with a classical heuristic that does not perform a search over the energy landscape or with a quantum heuristic that exploits tunneling.
In the present work we only showed a slight scaling advantage of the meta-heuristic over full-system SA and full-system PT. Demonstrating a stronger scaling advantage over a wider variety of competing heuristics remains an open task. If the present meta-heuristic could be modified to yield lower error, this may result in a stronger scaling advantage. This might be achieved by the following form of post-processing on the output of the present meta-heuristic: undecimate the spins in a randomly selected fragment, approximately optimize that fragment again, accept the new configuration only if it lowers the global energy, and repeat on a new random fragment. It would be straightforward to perform this post-processing in a way that preserves the linear scaling of the overall heuristic, but the scaling of the error in such a case would need to be empirically checked. Also, we note that one of the sources of error in our implementation of the meta-heuristic, finite temperature, could be eliminated without changing the time complexity: the ordinary tensors could be replaced by tropical tensors [68]. This would likely reduce the energy errors reported in this work at the same fragment sizes, however it would not be compatible with adapting the meta-heuristic to specialized photonic hardware in the manner discussed below.
In industrially-relevant combinatorial optimization problems, what is often desired is a diverse set of low-cost solutions, and some works have therefore addressed the problem of how to sample the low-energy configuration space without bias [69, 70, 71, 72]. We expect that our heuristic will not demonstrate good performance in this regard due to the locality of the single-fragment subroutine and also the enhanced sensitivity to finite numerical precision that arises from the exponential form of the Boltzmann weights that are intrinsic to our heuristic. This is consistent with the fact that in none of the tested problem instances was our heuristic able to obtain the ground state. This expectation is also consistent with the findings in the recent work of Ref. [73]. In that work, a branch-and-bound search strategy was combined with tensor-network-contraction marginal computations to approximately optimize quasi-two-dimensional classical spin lattices, and they found that the number of diverse solutions with less than 1% energy error that was generated with that method was a few orders of magnitude less than the number of such solutions obtained with other methods. For future work, therefore, we propose to hybridize the present method with other methods such that the present method is used to efficiently generate a few low-energy solutions that are in turn used as warm starts for other methods that have better performance in terms of exploring the low-energy configuration space, such as the algorithms in Refs. [70, 72]. It may be that such a hybridization produces a classical heuristic with an average-case time complexity that strongly challenges that of quantum heuristics for obtaining ground and low-energy states of very rugged energy landscape spin glasses.
Another augmentation of the method here could be along the lines of combining it with Monte Carlo similar to what is done in Refs. [74, 75, 76, 62]. In those works the sampling bias from approximate tensor-network contractions is corrected with the Metropolis scheme; in the case of the present method the Metropolis scheme could correct the bias that arises from finite fragment sizes.
We speculate that such improvements of the present algorithm may provide a new avenue for gaining insights into the properties of classical spin-glass models that are of interest in condensed matter. For example, questions remain open regarding the nature of the low-temperature phase of the cubic-lattice classical model [77]. Traditional Monte Carlo methods have been the state-of-the-art approach for this problem, but the hybrid methods that we propose would operate according to very different principles and may thereby yield new insights. As another example, Ref. [31] points out that the tunability of the tile-planted-solution model can allow for a systematic study of the interplay between disorder and frustration; the algorithms that we propose may provide a complementary route to traditional Monte Carlo methods in such a study as well.
The wall time of the present form of our meta-heuristic (i.e., using tensor-network contraction for the subsystem optimizations) does not compare favorably with full-system SA. However, we explained that the meta-heuristic may be parallelized through simultaneous use of multiple fragments. This would very substantially reduce the simulation time of our heuristic from what is reported here. Also, GPU acceleration has been shown to decrease the walltime of tensor-network contraction by over one order of magnitude [78]. Further, the computational cost of tensor-network contraction is dominated by matrix multiplications, for which specialized photonic hardware is under active development. Implementation of tensor-network contraction with such hardware could in principle yield substantially lower time and energy costs. However, it is unlikely that such hardware will be able to achieve beyond about 8 bits of numerical precision in the foreseeable future [79], whereas the results presented here were all with 64 bits of precision. Our (unshown) preliminary tests of our algorithm with float16 showed very poor results for the model and cubic-lattice tile planting model but modest results for Barahona’s two-level spin glass [2], which is a cubic-lattice reduction of Max-Cut on random three-regular graphs. Whether or not the present heuristic can be modified to generally yield sufficiently good warm starts when limited to 8-bit precision is an important open question.
It is possible that the heuristic presented in this work will not work well for short-range spin glasses that are reductions or embeddings of dense-graph combinatorial optimization problems. Besides, such optimization problems are more economically represented as dense-graph spin glasses. A tensor-network algorithm that can efficiently optimize dense-graph rugged-energy-landscape spin glasses is therefore desirable. Ref. [80] demonstrates that all-to-all coupled Ising spin glasses can sometimes be optimized by iterating over optimizations of randomly selected subsets of all-to-all coupled spins. In ongoing work using the Wishart planted ensemble [81], which is an all-to-all coupled classical Ising spin glass model with tunable ruggedness, we are investigating such an approach where the tensor-network machinery that we have demonstrated here is used to approximately optimize the random subsystems. As with the tensor-network-based approach that we demonstrated in the present work, such a tensor-network approach would be straightforward to generalize to beyond two-body interactions and beyond binary variables.
Acknowledgements.
We acknowledge discussions with Timothée Leleu, Sam Reifenstein, Victor Bastidas, Wangwei Lan, Johnnie Gray, Yu Tong, Tomislav Begušić, Garnet Chan, Salvatore Mandrà, and Sukhi Singh. We acknowledge Jack Raymond for pointing out Ref. [65] for the estimated ground state energy of the model in the thermodynamic limit. We acknowledge Humberto Munoz-Bauza for correspondence regarding the TAMC software package.Appendix A Simulated Annealing and Parallel Tempering details
Both simulated annealing and parallel tempering were implemented on Apple M2 Ultra CPU with 16 performance cores, 8 efficiency cores, and 128 GB of RAM. CPU-intensive background processes were terminated/disabled during the runs.
A.1 Simulated Annealing
We use the simulated annealing implementation in the dwave-samplers library [82]. Metropolis sweeps are performed according to a geometric temperature schedule. To minimize the TTS, the number of reads is kept fixed at 1 since we find this to always be sufficient to reach the target of interest for at least some choice of temperature range. For each problem instance we perform independent runs of SA for each possible temperature range that results from scanning the minimum between 0.1 and 2.4 in increments of 0.1 and the maximum between 3.0 and 13.0 in increments of 1. We choose the minimum TTS from these scans of for our reported data on all instances at all values of .
The chosen ranges of are determined through simulations on the ten instances of the class at : we find that the combination of minimum and maximum values of that optimizes the TTS on these instances always lies inside these respective ranges and not at their endpoints.
A.2 Parallel Tempering
We use the parallel tempering implementation in the TAMC software package [83]. We use a geometric temperature schedule and, for each problem instance, scan the minimum and maximum values of over the same values as for SA; we consider these ranges sufficient because in all of the ten instances of at the optimal minimum value of does not lie at the endpoints of the tested range, and in all but two the optimal maximum value does not lie at the endpoints of the tested range. By optimal we mean the value that minimizes the TTS. The number of values is scanned over 2, 3, and 4; we consider this a sufficient range since in all of the ten instances of at we find the optimal value to be either 2 or 3. Isoenergetic cluster moves are disabled. The number of sweeps is increased in increments of 1 starting from 20 to find the minimum time to reach the target for each possible combination of temperature range and number of values. This is done independently for each spin glass instance.
Appendix B Time-to-solution scaling data
Here we present data to verify that the actual time to solution of the tensor-network heuristic follows the theoretically-expected linear scaling.